[figure: 9.7 cm]

EFFECT OF Ni SUBSTITUTION ON THE MICROSTRUCTURE

AND MAGNETIC PROPERTIES IN FINEMET ALLOYS

J. Marcin^{a}, P. Švec^{b}, J. Kováč^{a} and I. Škorvánek^{a} aInstitute of Experimental Physics, Slovak Academy of Sciences, 043 53 Košice, Slovakia bInstitute of Physics, Slovak Academy of Sciences, 842 28 Bratislava, Slovakia

We have studied the effect of Ni substitution on the structural and magnetic properties of the Fe_{73.5−x}Ni_{x}Si_{13.5}Cu_{1}Nb_{3}B_{9} (x=0, 10, 20, 30) alloys. The increase of Ni concentration from 0 % to 30 %, causes a decrease of Curie temperature of amorphous matrix phase, T_{c}(am), from 593 K to 485 K. The results of the hysteresis loop measurements performed at room temperature have revealed a marked deterioration of soft magnetic properties with Ni substitution. A significant magnetic hardening was observed at the first stages of nanocrystallization for the sample containing 30 % Ni, while for the sample containing 10 % Ni it is visible only at advanced stages of crystallization. The effect of Ni substitution on the intergrain magnetic coupling was investigated via the temperature dependencies of coercivity in temperature range 300 K – 700 K. For all samples the coercivity markedly increases as temperature approaches T_{c}(am), indicating a decrease of exchange coupling between grains. This degradation of magnetic coupling at elevated temperatures is strongly enhanced by an increase of Ni content in the investigated samples.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : J. Marcin

Address for correspondence : Institute of Experimental Physics SAS Watsonova 43 043 53 Košice Slovakia

Email address : marcin@saske.sk